Data Span, Inc., automated its plant at the start of the current year and installed a flexible manufacturing system. The compRequired: 1-a. Compute the throughput time for each month. 1-b. Compute the delivery cycle time for each month. 1-c. ComputeRequired 1 Required 2 Required 3 1-a. Compute the throughput time for each month. 1-b. Compute the delivery cycle time for eaComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Evaluate the companys peRequired 1 Required 2 Required 3 3-a. (Month 5) Refer to the move time, process time, and so forth, given for month 4. Assume
Answer #1

Time Throughput Delivery Cycle Manufacturing time Cycle Efficiency (mee Month 1 8.3 days 33.3 days 27.7 Month 2 8.8 days 3612Delivery Cycle time for month 2 = (27.4 + 8.8 days i 3612 days Delivery Cycle time for month 3 =(3000 + 9.6) days = 39,6 daysThe throughput Time measure displays Unfavorable The Delivery cycle Time - days measure displays Unfavorable Manufacturing CyMonth 5 Month 6 Throughput Time 310 days 2.6 days Manufacturing Cycle Efficiency 66.7 % 76.9 % Calculation: 3a, wew Throughpu